OpenStudy (anonymous):

In triangle RST, angle S is a right angle and cscR=17/15. what is tan T?

OpenStudy (anonymous):



mathslover (mathslover):

Hello @berrypicking

mathslover (mathslover):

Well @vikrantg4 , your method has a great mistake. \[\large{\frac{RT}{TS} = \frac{17}{15}}\] is right but saying that RT = 17 and ST = 15 is wrong. as 2/4 = 1/2 but 2 is not equal to 1 and 4 is not equal to 2.

OpenStudy (anonymous):

I didn't say that but I do have the same diagram and am not sure what to do next

mathslover (mathslover):

now. @berrypicking . first of all : \[\large{\csc R = \frac{17}{15}}\] Right?

mathslover (mathslover):

and csc R = RT/TS , do you know why?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

no I know csc =1/sin

mathslover (mathslover):

Ok good. So now we have : \[\large{\frac{RT}{TS} = \frac{17}{15}}\] Mark this as equation (1) .

mathslover (mathslover):

see, as csc = 1/sin and sin = opposite / hypotenuse thus csc = hypotenuse / opposite opposite to angle R is ST and hypotenuse of the triangle is RT Thus csc R = RT/TS , got it?

OpenStudy (anonymous):

I know the opposiste and hypo but tell me why does cscR=RT/TS

mathslover (mathslover):

cosec = hypotenuse / opposite, agree?

OpenStudy (anonymous):

ok

mathslover (mathslover):

what is hypotenuse of the given triangle?

OpenStudy (anonymous):

RT

mathslover (mathslover):

right and opposite to angle R ?

OpenStudy (anonymous):

angle T

mathslover (mathslover):

Oh! We are talking about sides dude. So what is opposite "side" of angle R

OpenStudy (anonymous):

gotcha ST

mathslover (mathslover):

right so what is cosec R now?

OpenStudy (anonymous):

RT/ST

mathslover (mathslover):

Great work.

mathslover (mathslover):

So RT/ST = 17/15 , agreed?

OpenStudy (anonymous):

ok

mathslover (mathslover):

wait!

OpenStudy (anonymous):

so tan T = 15/RS. How do I find RS

mathslover (mathslover):

why tan T = 15/RS ?

OpenStudy (anonymous):

tan=opp/adj????

mathslover (mathslover):

yes. so?

OpenStudy (anonymous):

aren't I trying to find tan T what am I missing?

mathslover (mathslover):

see tan T = opp. / adj. is right. what is opposite side to angle T ?

OpenStudy (anonymous):

right to angle T -it is RS/ST

mathslover (mathslover):

opp. side to angle T is RS ... check that again and adjacent side to angle T is ST so tan T = RS/ST

mathslover (mathslover):

now , as tan T = \(\large{\frac{RS}{ST}}\) and from equation (1) \[\large{\frac{RT}{TS} = \frac{17}{15}}\] \[\large{\frac{RT}{17} \times 15 = TS}\] \[\large{\frac{15}{17} \times RT = TS}\] thus : \[\large{\tan T = \frac{RS}{ST} = \frac{RS}{\frac{15}{17}\times RT}}\] getting it?

mathslover (mathslover):

Now, \[\large{\tan T = \frac{RS}{RT} \times \frac{17}{15}}\] but RS/RT = cot R ( cot = adj. / opp. ) thus : \[\large{\tan T = \cot R \times \frac{17}{15}}\] Now our aim is to calculate cot R ... since : cot = 1 /tan thus cot R = 1 / tan R as tan R = sin R / cos R since sin R = 1/ cosec R and cosec R = 17/15 thus sin R = 15/17 since \(\large{\sin R = \sqrt{1- \cos^2 R}}\) Hence 15/17 = \(\sqrt{1- \cos^2 R}\) SQuare both sides: \[\large{\frac{225}{289} = 1 - cos^2 R}\] \[\large{1- \frac{225}{289} = \cos ^2 R}\] \[\large{ \frac{64}{289} = \cos^2 R}\] Now we have cos R = 8/17 hence tan R = sin R / cos R = (15/17)/(8/17) = 15/8 got it?